### v1

@@ 3.2 Preprocessing
Preprocessing plays a pivotal role in reducing noise and standardizing inputs. A recent study [49] investigates the impact of popular preprocessing methods on transformer models. The findings indicate that, in some cases, with an appropriate preprocessing strategy, even a simple Naïve Bayes classifier can outperform the best-performing transformer by 2% in accuracy. Other studies provide direct evidence that certain preprocessing techniques enhance the effectiveness of hate and offensive speech detection [22] as well as toxicity detection [35].
@@ 3.2.1 Data Cleaning
Effective data cleaning is essential for preprocessing text in toxicity detection tasks, as extraneous and misleading content can interfere with model performance. One common step is URL removal, as URLs often introduce external references that may skew predictions. URLs are typically detected using regex patterns that match common formats such as ”http://” or domain names (e.g., www.example.com). Removing these links prevents models from relying on unrelated external content, ensuring they focus solely on textual features. This technique has been widely applied in prior studies [26, 6] to reduce biases. However, its overall effectiveness remains a topic of discussion, as studies have shown that URL removal has a negligible impact on sentiment classification accuracy and is often outperformed by more advanced preprocessing techniques [27].
@@
Another important step is contraction expansion, where informal contractions like ”can’t” or ”won’t” are expanded into their full forms (”cannot” and ”will not”). This transformation standardizes text, making it easier for models to parse words correctly and recognize them as distinct entities. Studies such as Jianqiang and Xiaolin [27] have shown that contraction expansion improves accuracy and F1-scores in sentiment classification, demonstrating its relevance in text preprocessing pipelines. However, its effectiveness can vary across domains, as expanding contractions may introduce unnecessary verbosity or alter the tone of informal text, potentially impacting sentiment analysis in conversational contexts [1].
@@
Symbol removal is another preprocessing step that enhances text clarity. Symbols such as ”@”, ”#”, or excessive punctuation (e.g., ”!!!”) are common in social media text but do not necessarily contribute meaningful information for toxicity detection. Their presence can introduce noise and cause models to overfit to irrelevant patterns. Removing unnecessary symbols [3, 52] helps streamline the data and ensures that predictions are based on meaningful language rather than stylistic artifacts. Similarly, repetition elimination addresses exaggerated expressions commonly found in informal text, where characters (e.g., ”loooove” → ”love”) or words (e.g., ”really really good”) are repeated. If not addressed, models may treat these variations as distinct features, leading to unnecessary complexity. By simplifying repeated elements [8], we maintain the intended meaning while improving model consistency. However, it’s important to note that while symbol removal can reduce noise, it may also strip away context that is crucial for understanding the tone or intent behind a message. For instance, emoticons often convey contextual meaning, which can be valuable for sentiment analysis [1]. Thus, indiscriminate removal of symbols may risk losing important nuances in informal text, particularly on platforms where symbols play a significant role in communication.
@@
For domain-specific datasets, particularly those containing programming-related text, additional preprocessing steps are necessary to distinguish between toxic language and structural code components. One such technique is identifier splitting, which breaks down compound variable names into separate tokens. Many programming identifiers, such as ”isToxic”, use camelCase or snake_case formatting, making them difficult for models trained on natural language text to interpret. Splitting these identifiers into distinct words (”is” and ”toxic”) allows models to focus on meaningful linguistic elements rather than code syntax, enhancing their ability to detect toxicity [29].
@@
Another preprocessing method specific to code-related datasets is programming keyword removal. Programming languages contain syntax elements such as keywords (”if”, ”else”, ”return”), operators (”+”, ”=”, ”&&”), and delimiters (”{”, ”}”, ”;”), which are crucial for code execution but generally irrelevant for toxicity detection. By removing these elements, models can concentrate on meaningful text components, such as comments or string literals, where toxicity is more likely to be present. This approach has been validated in studies like HaCohen-Kerner et al. [23], which identified the removal of HTML objects and syntax elements as a key step in effective preprocessing. Filtering out programming-specific tokens helps models assess toxicity more accurately by preventing structural code artifacts from influencing the analysis.
@@ 3.2.2 Feature Engineering
Feature engineering plays a crucial role in improving model performance by incorporating customized attributes that enhance the detection of toxic content. One key technique is adversarial pattern identification, which focuses on detecting and mitigating adversarial inputs designed to bypass toxicity filters. These inputs often include obfuscated slurs or profanities crafted to evade automated detection. For instance, Hartvigsen et al. [24] employ adversarial pattern generation to identify implicit toxicity, improving the robustness of toxicity classifiers against evasive language tactics. However, this approach can inadvertently perpetuate bias by focusing on specific, often marginalized, groups targeted by toxic language. It risks amplifying the very patterns it aims to detect, rather than addressing the root causes of toxic discourse. For example, a study [43] demonstrates that toxicity detection models trained on existing datasets often exhibit significant disparities in performance across different demographic groups, with higher false positive rates for language associated with minority identities. This indicates that focusing narrowly on adversarial pattern identification, without considering broader social context and potential biases, can exacerbate existing inequities in toxicity detection.
@@
Another widely used approach is profane word counting, which involves calculating the density of harmful or offensive words in a given comment. This method provides a numerical representation of potential toxicity, serving as an important feature for classifier training. Research by Bhat et al. [10] outlines a taxonomy for workplace toxicity and highlights profanity as a key indicator of offensive language. Similarly, d’Sa et al. [17] include profane word counting in their preprocessing pipeline to enhance classification accuracy.
@@
Embedding features leverage word or sentence embeddings, such as Word2Vec, GloVe [60], and BERT [12], to represent text in a high-dimensional continuous space. These embeddings capture semantic relationships between words, allowing models to better understand nuanced language patterns associated with toxicity. By transforming text into a structured numerical format, embedding features enhance the model’s ability to detect implicit and context-dependent toxic language. Embedding features, while powerful for capturing semantic relationships, risk amplifying societal biases present in their training data. Studies like Blodgett et al. [11] reveal that such biases can lead to disproportionate toxicity flagging of language associated with marginalized groups, necessitating careful mitigation strategies.
@@
Finally, contextual features are crucial for analyzing user interactions and conversational history to disambiguate toxicity. Isolated toxic statements may appear ambiguous without considering the surrounding discussion. For example, Pavlopoulos et al. [39] investigate the importance of conversational context in toxicity classification, demonstrating that models incorporating thread structure and user history outperform those relying solely on individual comments. By integrating contextual features, models can differentiate between harmful intent and benign language, improving overall classification accuracy. However, processing long conversational histories introduces a significant computational burden, hindering real-time applications due to the memory and processing demands of capturing long-range dependencies. Research into transformer efficiency [53] highlights the ongoing efforts to address this challenge.
@@ 3.2.3 Sampling techniques
To ensure a balanced representation of toxic and nontoxic samples, some studies [33, 46] employ stratified sampling based on external toxicity scores, such as those from Google’s Perspective API. This approach prevents bias by balancing the distribution of toxicity likelihoods across subsets and ensures that the model is exposed to a diverse range of samples, avoiding overfitting to one class. However, while stratified sampling aims for representativeness based on pre-defined criteria, it may not necessarily select the most informative samples for model training. As an alternative,  active learning is a technique that strategically selects the most valuable data points for labeling, focusing on those that will maximize the model’s learning with minimal labeled data. It can achieve comparable or even better performance with significantly fewer labeled samples by strategically selecting those that maximize model learning [19]. Therefore, active learning could be explored as an alternative or complementary technique to stratified sampling, especially when the cost of labeling is a concern.
@@ 3.3 Detection & Mitigation
The detection and mitigation of toxicity in SE communication have evolved significantly, transitioning from simple, keyword-based approaches to sophisticated, context-sensitive models. Current strategies employ both general-purpose and domain-specific language models, reflecting an essential balance between broad applicability and specialized performance. Recent advances highlight the necessity of tailoring detection systems to SE-specific language, given the domain’s unique terminologies and nuanced communication patterns. The following sections detail these distinctions and explore various architectural designs and task-specific approaches in toxicity management within the SE domain.
@@
Large pre-trained language models, such as GPT-3, GPT-4, and Gemini, demonstrate notable flexibility in toxicity detection tasks due to their zero-shot and few-shot capabilities. Mishra and Chatterjee [34] showcased how ChatGPT effectively identified toxic comments on GitHub without requiring domain-specific fine-tuning, leveraging its training on diverse datasets. Similarly, Sultana and Begum Kali [50] demonstrated ChatGPT’s success in detecting gender biases and stereotypes, such as maternal insults and subtle stereotyping, in SE communications. Generative pre-training enables these models to adapt to a range of tasks with minimal input. However, performance evaluations reveal limitations: Shafikuzzaman et al. [48] showed that while GPT-4o performs reasonably well in smaller SE datasets, it falls behind fine-tuned models like seBERT when applied to larger datasets such as GIT and ARC. This highlights a critical trade-off between flexibility and performance when domain-specific nuances are involved.
@@
Focusing on the specific requirements of SE discussions, domain-specific models fine-tuned on SE datasets consistently outperform general-purpose models in both accuracy and contextual relevance. Sarker [45] emphasized the importance of pre-processing SE-specific language, noting that general-purpose models often misclassify technical terms such as ”kill” or ”junk” as toxic. The other benchmark study [46] demonstrated how retraining in SE-specific data significantly improves performance. For example, experiments that evaluated tools like DPCNN and BFS after retraining on SE datasets such as Code Review and Gitter Ethereum revealed considerable improvements in key metrics such as precision, recall, and F score. The accompanying Figure 3 and Figure 4 highlight these findings, with F scores for DPCNN and BFS reaching 0.406 and 0.366, respectively, on Code Review, and 0.652 and 0.660 on Gitter Ethereum. These metrics underscore the value of retraining models to handle SE-specific language more effectively.
@@
In addition, Sarker et al. [47] introduced a span-based toxicity detector that identifies specific toxic phrases in the SE text, providing improved interpretability for moderators. Beyond binary classification, Cheriyan et al. [13] analyzed toxicity across platforms like GitHub and Stack Overflow, advocating for SE-specific context-aware solutions to reduce conflicts and enhance community dynamics. Recent benchmarks also highlight seBERT’s macro F1-scores exceeding 90% on datasets like GIT, reflecting its ability to capture domain-specific nuances with unparalleled precision [48]. These examples emphasize that models tailored to the SE domain are essential for achieving superior performance in detecting and mitigating toxicity.
@@
Figure 3: Code Review Dataset (Comparing Methods Before & After Retrain on SE Dataset)
@@
Figure 4: Gitter Ethereum Dataset (Comparing Methods Before & After Retrain on SE Dataset)

### v2

@@ 3.2 Pre-processing
Pre-processing plays a pivotal role in reducing noise and standardizing inputs. A recent study [siino2024text] investigates the impact of popular pre-processing methods on transformer models. The findings indicate that, in some cases, with an appropriate pre-processing strategy, even a simple Naïve Bayes classifier can outperform the best-performing transformer by 2% in accuracy. Other studies provide direct evidence that certain pre-processing techniques enhance the effectiveness of hate and offensive speech detection [glazkova2023comparison] as well as toxicity detection [mohammad2018preprocessing].
@@ 3.2.1 Data cleaning
Effective data cleaning is essential for pre-processing text in toxicity detection tasks, as extraneous and misleading content can interfere with model performance. One common step is URL removal, as URLs often introduce external references that may skew predictions. URLs are typically detected using regex patterns that match common formats such as “http://” or domain names (e.g., “www.example.com”). Removing these links prevents models from relying on unrelated external content, ensuring they focus solely on textual features. This technique has been widely applied in prior studies [jemai2021sentiment, athar2021sentimental] to reduce biases. However, its overall effectiveness remains a topic of discussion. Studies by jianqiang2017comparison, for instance, observe that URL removal often has a negligible impact on sentiment classification accuracy, with fluctuations in accuracy and F1-measure generally limited to +/-0.5% and, in many cases, no change at all. This suggests that it is often outperformed by more advanced pre-processing techniques.
@@
Another important step is contraction expansion, where informal contractions like “can’t” or “won’t” are expanded into their full forms (“cannot” and “will not”). This transformation standardizes text, making it easier for models to parse words correctly and recognize them as distinct entities. Studies have shown that contraction expansion improves accuracy and F1-scores in sentiment classification, demonstrating its relevance in text pre-processing pipelines. For instance, jianqiang2017comparison report accuracy gains of up to 6.08% and F1-score improvements of up to 6.85% on certain datasets and classifiers after expanding acronyms. However, its effectiveness can vary across domains, as expanding contractions may introduce unnecessary verbosity or alter the tone of informal text, potentially impacting sentiment analysis in conversational contexts [aadil2023exploring].
@@
Symbol removal is another pre-processing step that enhances text clarity. Symbols such as “@”, “#”, or excessive punctuation (e.g., “!!!”) are common in social media text but do not necessarily contribute meaningful information for toxicity detection. Their presence can introduce noise and cause models to overfit to irrelevant patterns. Removing unnecessary symbols [anbukkarasi2020analyzing, tan2023survey] helps streamline the data and ensures that predictions are based on meaningful language rather than stylistic artifacts. Similarly, repetition elimination addresses exaggerated expressions commonly found in informal text, where characters (e.g., “loooove” → “love”) or words (e.g., “really really good”) are repeated. If not addressed, models may treat these variations as distinct features, leading to unnecessary complexity. By simplifying repeated elements [aziz2020twitter], we maintain the intended meaning while improving model consistency.
However, it’s important to note that while symbol removal can reduce noise, it may also strip away context that is crucial for understanding the tone or intent behind a message. For instance, emoticons often convey contextual meaning, which can be valuable for sentiment analysis [aadil2023exploring]. Thus, indiscriminate removal of symbols may risk losing important nuances in informal text, particularly on platforms where symbols play a significant role in communication. This nuanced impact on model performance is evident from studies like jianqiang2017comparison, which reports that while some models gained up to 12.52% in F1-score after reverting repetitions, others experienced significant losses, such as a 12.11% accuracy decrease for Naive Bayes on certain datasets.
@@
For domain-specific datasets, particularly those containing programming-related text, additional pre-processing steps are necessary to distinguish between toxic language and structural code components. One such technique is identifier splitting, which breaks down compound variable names into separate tokens. Many programming identifiers, such as “isToxic”, use camelCase or snake_case formatting, making them difficult for models trained on natural language text to interpret. Splitting these identifiers into distinct words (“is” and “toxic”) allows models to focus on meaningful linguistic elements rather than code syntax, enhancing their ability to detect toxicity. This approach has shown positive impacts on model performance; in the context of code similarity detection, karnalim2020preprocessing report F-score increases of around 14% and recall improvements of around 35% on Python datasets after applying identifier removal.
@@
Another pre-processing method specific to code-related datasets is programming keyword removal. Programming languages contain syntax elements such as keywords (“if”, “else”, “return”), operators (“+”, “=”, “&&”), and delimiters (“{”, “}”, “;”), which are crucial for code execution but generally irrelevant for toxicity detection. By removing these elements, models can concentrate on meaningful text components, such as comments or string literals, where toxicity is more likely to be present. This approach has been validated in studies like hacohen2020influence, which identified the removal of HTML objects and syntax elements as a key step in effective pre-processing. Filtering out programming-specific tokens helps models assess toxicity more accurately by preventing structural code artifacts from influencing the analysis.
To guide practitioners in applying these pre-processing techniques, we propose a sequenced workflow, particularly for toxicity detection in software engineering contexts. A general pipeline would typically involve: 1) removing URLs, 2) expanding contractions, 3) eliminating repetitions, and 4) performing symbol removal. For domain-specific datasets, such as those containing programming-related text, additional steps would then be applied: 5) splitting identifiers, and 6) filtering out programming keywords. The precise order and inclusion of steps may vary based on dataset characteristics and model requirements, but this general sequence provides a robust starting point for effective data cleaning.
@@ 3.2.2 Feature engineering
Feature engineering plays a crucial role in improving model performance by incorporating customized attributes that enhance the detection of toxic content. One key technique is adversarial pattern identification, focusing on detecting and mitigating inputs designed to bypass toxicity filters, such as obfuscated slurs or profanities using patterns like f[\\*@_\\-]+ck for “fuck” or s#+it for “shit” that account for character substitutions. For instance, hartvigsen2022toxigen employ adversarial pattern generation to identify implicit toxicity, enhancing classifier robustness against evasive language. However, this approach risks inadvertently perpetuating bias by focusing on specific, often marginalized, groups targeted by toxic language, potentially amplifying the very patterns it aims to detect rather than addressing root causes. A study [sap2019risk] further demonstrates that toxicity detection models trained on existing datasets often exhibit significant performance disparities across demographic groups, with higher false positive rates for language associated with minority identities. This indicates that a narrow focus on adversarial pattern identification, without considering broader social context and potential biases, can exacerbate existing inequities in toxicity detection.
@@
Profane word counting is another widely used approach, which calculates the density of harmful or offensive words in a comment, typically as the ratio of identified profane words to the total word count. This numerical feature serves as a quantifiable indicator of potential toxicity for classifier training. For example, research by bhat2021say outlines a taxonomy for workplace toxicity and highlight profanity as a key indicator of offensive language. Similarly, d2019towards include profane word counting in their pre-processing pipeline to enhance classification accuracy.
@@
Embedding features leverage word or sentence embeddings, such as Word2Vec, GloVe [zhang2019hate], and BERT [caselli2020hatebert], to represent text in a high-dimensional continuous space. These embeddings capture semantic relationships between words, allowing models to better understand nuanced language patterns associated with toxicity. By transforming text into a structured numerical format, embedding features enhance the model’s ability to detect implicit and context-dependent toxic language. Embedding features, while powerful for capturing semantic relationships, risk amplifying societal biases present in their training data. Studies like blodgett2020language reveal that such biases can lead to disproportionate toxicity flagging of language associated with marginalized groups, necessitating careful mitigation strategies.
@@
Finally, contextual features are crucial for analyzing user interactions and conversational history to disambiguate toxicity. Isolated toxic statements may appear ambiguous without considering the surrounding discussion. For example, pavlopoulos2020toxicity investigate the importance of conversational context in toxicity classification, demonstrating that models incorporating thread structure and user history outperform those relying solely on individual comments. By integrating contextual features, models can differentiate between harmful intent and benign language, improving overall classification accuracy. However, processing long conversational histories introduces a significant computational burden, hindering real-time applications due to the memory and processing demands of capturing long-range dependencies. Research into transformer efficiency [tay2022efficient] highlights the ongoing efforts to address this challenge.
Beyond the generation of effective features, careful feature selection and dimensionality reduction are critical steps to prevent overfitting, reduce computational complexity, and enhance model interpretability. Techniques such as mutual information [battiti1994using] and chi-square tests [yang1997comparative] can be employed to evaluate the statistical dependence between features and the target variable, thereby identifying the most predictive attributes. Additionally, using L1-regularized models (e.g., Lasso regression [tibshirani1996regression]) inherently performs feature selection by shrinking the coefficients of less important features to zero. Implementing these strategies guides practitioners in curating a lean yet powerful feature set, optimizing model performance and generalization capabilities.
@@ 3.2.3 Sampling techniques
To ensure balanced representation of toxic and non-toxic samples, studies often employ stratified sampling based on external toxicity scores, such as those from Google’s Perspective API [miller2022did, sarker2020benchmark]. Given the rarity of toxic content, practical guidelines involve stratifying samples into specific toxicity score ranges; for instance, sarker2020benchmark used 0.1-interval groups for scores below a 0.5 threshold to ensure diverse representation across the non-toxic spectrum. This approach balances the distribution of toxicity likelihoods across subsets, exposing the model to diverse samples and preventing overfitting to a single class. However, while stratified sampling prioritizes representativeness, it may not always select the most informative samples for model training. To tackle data imbalance, other over-sampling techniques like SMOTE [chawla2002smote] can be utilized.
As an alternative, active learning strategically selects the most valuable data points for labeling, aiming to maximize model learning with minimal labeled data. For instance, figueroa2012active demonstrated that active learning can achieve 90% accuracy with approximately 33% fewer instances than random sampling on certain datasets. Beyond just efficiency, active learning also offers an advantage in addressing data imbalance, as it focuses on efficiently acquiring informative samples from the minority class. Consequently, active learning could serve as an effective alternative or complementary technique to stratified sampling, particularly when labeling costs are a significant consideration.
@@ 3.3 Detection & Mitigation
The detection and mitigation of toxicity in SE communication have evolved dramatically. It has moved from simple, keyword-based methods to sophisticated, context-sensitive models. Current strategies balance the broad domain coverage of general-purpose models with the higher precision of SE-tuned, domain-specific ones. Given the unique terminology and nuanced patterns of SE text, recent work underscores the need for SE-specific pre-processing and fine-tuning. In the sections below, we first compare general-purpose and SE-specific approaches, then examine different architectures and task-level methods for toxicity management in SE.
@@
Large pre-trained language models, such as GPT-3, GPT-4, and Gemini, demonstrate notable flexibility in toxicity detection tasks due to their zero-shot and few-shot capabilities. mishra2024exploring showcased how ChatGPT effectively identified toxic comments on GitHub without requiring domain-specific fine-tuning, leveraging its training on diverse datasets. Similarly, sultana2024exploring demonstrated ChatGPT’s success in detecting gender biases and stereotypes, such as maternal insults and subtle stereotyping, in SE communications. Generative pre-training enables these models to adapt to a range of tasks with minimal input. However, performance evaluations reveal limitations: shafikuzzaman2024empirical showed that while GPT-4o performs reasonably well in smaller SE datasets, it falls behind fine-tuned models like seBERT when applied to larger datasets such as GIT and ARC. This highlights a critical trade-off between flexibility and performance when domain-specific nuances are involved.
@@
Focusing on the specific requirements of SE discussions, domain-specific models fine-tuned on SE datasets consistently outperform general-purpose models in both accuracy and contextual relevance. sarker2022built emphasized the importance of pre-processing SE-specific language, noting that general-purpose models often misclassify technical terms such as “kill” or “junk” as toxic. The other benchmark study demonstrated how retraining in SE-specific data significantly improves performance. For example, sarker2020benchmark’s experiments evalutated tools like DPCNN and BFS after retraining on SE datasets such as Code Review and Gitter Ethereum revealed considerable improvements in key metrics such as precision, recall, and F score. The accompanying Figure 3 and Figure 4 highlight these findings, with F scores for DPCNN and BFS reaching 0.406 and 0.366, respectively, on Code Review, and 0.652 and 0.660 on Gitter Ethereum. These metrics underscore the value of retraining models to handle SE-specific language more effectively.
@@
In addition, sarker2023toxispanse introduced a span-based toxicity detector that identifies specific toxic phrases in the SE text, providing improved interpretability for moderators. Beyond binary classification, cheriyan2021towards analyzed toxicity across platforms like GitHub and Stack Overflow, advocating for SE-specific context-aware solutions to reduce conflicts and enhance community dynamics. Recent benchmarks also highlight seBERT’s macro F1-scores exceeding 90% on datasets like GIT, reflecting its ability to capture domain-specific nuances with unparalleled precision [shafikuzzaman2024empirical]. These examples emphasize that models tailored to the SE domain are essential for achieving superior performance in detecting and mitigating toxicity.
@@
Figure 3: Code review dataset (comparing methods before & after retrained on SE Dataset)
@@
Figure 4: Gitter Ethereum dataset (comparing methods before & after retrained on SE dataset)
Table 8: Cloud LLM APIs and corresponding operational costs
Model Provider Input (\$/M tokens) Output (\$/M tokens) Latency (p99, s)
GPT-4.1 OpenAI 2.00 8.00 2.66
Gemini 1.5 Flash Google 0.60 2.40 0.92
Table 8 summarizes the prices and latencies of the commercial LLM APIs used in our experiments in Section 5, reflecting the conditions at the time the experiments were conducted [openai2024gpt41, Google2025VertexAIPricing]. For the latest and most accurate information regarding model availability and pricing, please refer to the providers’ official websites, as these models are frequently updated and iterated. Because these services are pay-as-you-go, teams can simply pick the SKU that matches their quality target and budget. Take GPT-4.1 as an example: suppose we need to moderate 60 comments every hour, each averaging 50 English words. OpenAI’s tokenizer turns English prose into roughly one token for every 0.75 word, so a 50-word comment is about 67 tokens. Multiplying by 60 comments yields around 4,000 input tokens per hour. If we budget a short, one-token label plus a few JSON characters—say 5 output tokens per comment—that adds another 300 output tokens. The total comes to about \$0.01 per hour.
A second route is to deploy the model to cloud service providers like Amazon Web Service. An inexpensive yet GPU-equipped choice is g4dn.xlarge (1 $\times$ NVIDIA T4, 16 GB) [AWS2025G4dnXlarge]. The on-demand price in us-east-1 is \$0.526 per hour. Adding a 30 GB gp3 EBS volume for the OS and model weights costs about \$2.40 per month at the published \$0.08 GB-month rate. Network pricing is favourable: inbound traffic is free and the first 100 GB of data that leaves AWS each month is also free; only excess egress is billed (starting near \$0.09 GB). With seBERT’s $\sim$1 ms inference time per 128-token comment, one g4dn.xlarge can serve roughly one million requests in 20 minutes, so an eight-hour moderation shift adds about \$4 of compute.
The advantages and disadvantages of each option are presented in Table 9.
Table 9: Comparison of LLM deployment choices
Deployment Choice Advantages Disadvantages
Commercial LLM API (e.g., GPT-4.1) Zero ops, instant scalability. Pay-as-you-go pricing (as low as \$0.01/h for the 60-comment example). No GPU provisioning required. Latency depends on network conditions. Costs grow linearly with usage (per-token charges). Data leaves your VPC.
Cloud Self-Host (EC2 g4dn.xlarge) Fixed hourly rate (\$0.526/h) with full GPU control. Domain-tuned models (e.g., seBERT) may achieve higher F1 at similar runtime cost. 100 GB outbound traffic/month is free. Charges apply even when idle (unless instance stopped). Requires manual patching, monitoring, and scaling. Additional EBS storage at \$0.08/GB-mo.
On-Premise Hardware Data never leaves local network. No recurring cloud fees after purchase. Sub-millisecond latency possible with PCIe GPUs. High upfront cost. Depreciation, power, and cooling overhead. Full responsibility for redundancy, security, and capacity planning.